# Dynamic survival analysis: modelling the hazard function via ordinary differential equations

### v3

@@ 3 Examples of hazard models defined through ODEs
In this section we present two examples that we consider of practical interest. The first example shows that the ODE formulation (1) can be used to represent common probability distributions with positive support, but also contains extensions, including distributions with a hazard function $h$ with positive and finite initial condition $h(0)$. A particular model, obtained by specifying the hazard function as the solution to the Logistic growth model (Boyce et al., 2021), is presented. This model has been largely used in ecology for modelling population growth, and represents a classical example in most textbooks on ODEs. This model does not include hidden state variables $q$ and has an analytic solution. The second example corresponds to defining the hazard function as the solution of a version of the competitive Lotka-Volterra model (Kot, 2001; Murray, 2002). The competitive Lotka-Volterra model is often used to model species in competition, but here we provide an alternative use and interpretation of this model to represent the interaction of the hazard associated with a disease and the response coming from a combination of therapeutic interventions and the immune system (at the population level). This model, presented in Section 3.2, has a hidden state variable $q$ and no analytic solution, so a numerical solver is employed to approximate its solution.
@@ 3.1 Probability distributions defined through a first order ODE
The following remark shows that our formulation (1) can be used to represent common absolutely continuous probability distributions with positive support by specifying a Bernoulli-type differential equation (Boyce et al., 2021).
@@
The term $a(t)$ in (4) may explicitly depend on $t$, which implies that the corresponding ODE is non-autonomous. If the term $a(t)$ in (4) can be written in terms of $h(t)$ only, then the corresponding ODE is autonomous. Thus, the term $a(t)$ in (4) can be interpreted as an “autonomy coefficient”. For example, the Weibull hazard function, $h(t)=\beta\kappa t^{\kappa-1}$, may be obtained as the solution to the ODE (4) with $a(t)=(\kappa-1)\left(\dfrac{\beta\kappa}{h(t)}\right)^{\frac{1}{\kappa-1}}-h(t)$, for $\kappa\neq 1$. This is an autonomous ODE. In contrast, for the log-normal distribution $a(t)=\frac{\mu-\sigma^{2}-\log(t)}{\sigma^{2}t}$, which implies that the ODE (4) associated with the log-normal distribution is non-autonomous. This in line with the classical result that, under mild regularity conditions, autonomous scalar ODEs have monotonic solutions (Wallach, 1948). These ideas are summarised in our context in the following remark.
@@
$h^{\prime}(t)=\psi_{\theta}(h),\quad h(t_{0})=h_{0},$
where $\theta\in\Theta$, $\psi_{\theta}$ is a continuous function, and $h(t)>0$ is the corresponding solution. Then, $h(t)$ is monotonic. Moreover, Let $h(t)>0$ be the solution to a first order (one-dimensional) ODE
$h^{\prime}(t)=\psi_{\theta}(h,t),\quad h(t_{0})=h_{0},$ (5)
If $h(t)$ is monotonic, then the ODE (5) is autonomous. That is, $\psi_{\theta}(h,t)$ depends only on $h$.
@@
Some references have claimed that equations of type (4) can be seen as a characterisation of absolutely continuous distributions with positive support (see Baran and Coughlin, 1987 and Onose, 1986). However, our formulation (1) indicates otherwise, as formulations with hidden state variables cannot be reduced to a Bernoulli ODE of type (4). Moreover, for most of the commonly used distributions, the corresponding hazard function $h(t)$ is the solution to an equation of type (4), with $\lim_{t\to 0}h(t)$ restricted to be either 00 or $\infty$. Next, we present an example of (1) without hidden state variables, and with positive and finite initial conditions on $h$.
3.1.1 Logistic growth hazard model
The logistic growth hazard model is defined through the following system of ODEs,
@@ 3.1.1 Logistic growth hazard model
where $\lambda>0$ represents the intrinsic growth rate of the hazard function, $\kappa>0$ represents the upper or lower bound of the hazard function, and $h_{0}>0$ is the value of the hazard function at $t=0$. The logistic growth ODE is an autonomous Bernoulli ODE of type (4), with $a(t)=\lambda-(1+\frac{\lambda}{\kappa})h(t)$ in (4). This ODE has the following analytic solution
@@
is a realisation of the logistic growth model. An advantage of model (6) is that it does not force the hazard function to be either 00 or $\infty$ at $t=0$, in contrast to the Weibull distribution (among other distributions). Moreover, the shapes of the hazard function are also increasing or decreasing. The parameters have a clear interpretation as $h_{0}$ represents the value of the hazard function at $t=0$, $\kappa$ represents the asymptotic point of stability of the hazard function, and $\lambda$ represents the growth rate of the hazard function. Figure 1 shows some examples of the logistic growth hazard model. Only for comparisons, since there exists and analytic solution. These were also calculated using the “Livermore Solver for Ordinary Differential Equations with Automatic method switching” (LSODA, SciPy implementation). The analytic solution is plotted for $h_{0}=\kappa/10$, for comparison. In all cases, we can notice how $h(t)\rightarrow\kappa$ as $t\to\infty$.
@@ 3.2 Hazard-Response model
The qualitative analysis of this system is more complicated than that of the simpler logistic-growth model studied in Section 3.1.1, and may be found in Murray (2002), section 3.5. Briefly, this model has four steady state solutions. These steady states can be found by solving the equation $\psi_{\bm{\theta}}(h,q)=\mathbf{0}$, which defines four cases: Case 1, $h=0,q=0$; Cases 2 and 3, $h=\kappa,q=0$ and $h=0,q=\kappa$, respectively; and Case 4, $h=h^{*},q=q^{*}$, if both positive, where
@@ 4 Inference
where we are including the initial conditions ${\bf Y}_{0}$ as unknown parameters in this equation. In some cases, as we will discuss in the simulation study and the applications, in Sections 5-6, it is possible to accurately estimate the initial conditions. However, in other cases, as discussed in Section 6.2, the data contains little or no information about these parameters, leading to practical non-identifiability, which is reflected as flat likelihood surfaces (Cole, 2020). Indeed, the initial conditions are commonly fixed at reasonable values in many practical applications of ODE models (Simpson et al., 2022). Once a solution (analytic or numerical) to the system of ODEs (1) has been obtained, we can evaluate this likelihood function at specific values of the parameters $\bm{\theta}$. In particular, for systems of ODEs without analytic solution and once a solver has been specified, we can retrieve the hazard and cumulative hazard functions, $h(t\mid\bm{\theta},{\bf Y}_{0})$ and $H(t\mid\bm{\theta},{\bf Y}_{0})$, at the sequence of observed time points ${\bf t}=\{t_{1},\dots,t_{n}\}$. This allows for calculating the maximum likelihood estimates (MLEs) of the parameters $\bm{\theta}$ and the initial conditions ${\bf Y}_{0}$ using general-purpose optimisation algorithms (e.g. optim or nlminb in R).
Although it is not possible to come up general prior choices for the parameters of the hazard models obtained via (1), as different models contain different types of parameters, the ease of interpretation of those parameters facilitates choosing either informative or weakly informative priors. The use of improper priors would require a case-by-case analysis to check the propriety of the posterior distribution.
@@ 5.1 Simulation scenarios
For simulation scenario 1, we simulate $M=250$ samples of sizes $n=250,500,1000,5000$ from the logistic growth hazard model (6) with parameters $(\lambda,\kappa,h_{0})=(0.5,0.05,3.5)$. These parameter values produce a decreasing hazard function, which starts at $h_{0}=3.5$, and decreases at a relatively fast rate, $\lambda=0.5$, to the carrying capacity value $\kappa=0.05$. These samples are simulated using the procedure described in Section 3.1.1. Administrative censoring times (i.e. fixed) are used to induce censoring rates of approximately $50\%$ and $25\%$.
@@ 5.2 Results
Tables 34 show the summaries of the results for simulation scenario 2. We can see that the Bayes estimates are close to the true values of the parameters, but exhibit some degree of bias which does not disappear with increasing sample size. This seems to be a result of the step size used in the simulation and the interval length. In order to reduce this bias, it is required to use a smaller step size and larger interval length. On the other hand, we can see that this configuration allows one to recover the shape of the hazard with high accuracy, as shown in Figure 7, which presents the true hazard function together with the $95\%$ predictive intervals.
@@ 6.1 Case study I: Leukemia data
In this application, we analyse the LeukSurv data set from the spaBayesSurv R package, using the Logistic hazard model (6) presented in Section 3.1.1. This data set contains information about the survival of $n=1043$ patients with acute myeloid leukemia. The maximum follow-up time was $13.6$ years for this data set. By the end of the follow-up $879$ individuals died and $164$ were still alive.
We simulate from the posterior distribution of the parameters $(\lambda,\kappa,h_{0})$ of the Logistic hazard model (6) using the t-walk sampler (Christen and Fox, 2010). The t-walk is a self-adaptive MCMC algorithm that requires no tuning parameters. We obtain $500,000$ posterior samples, and use a burn-in period of $20,000$ iterations and a thinning period of $167$ iterations. This is the estimated integrated auto-correlation time (IAT, see details of output analysis in Molina-Muñoz and Christen, 2022), and divided by the dimension (here, three) results in $55.6$, leading to an effective sample size of $480,000/167\approx 2870$. Figure 2 shows the corresponding posterior predictive hazard and survival functions for this population of cancer patients. The posterior predictive hazard function starts at $h_{0}$ (posterior mean $\approx 3.6$) and sharply decreases to $\kappa$ ($95\%$ posterior credible interval $(0.012,0.125)$) within the first 4 years. These results are in line with the progression of this type of cancer tumours, which tend to develop quickly with only a small portion of the population responding positively to treatments (Shallis et al., 2019), leading to poor prognosis and survival. Now, comparing the Logistic hazard model against the Weibull distribution using the Bayesian information criterion (BIC): the Logistic hazard model has a $BIC=1862.6$, while fitting a Weibull distribution to the same data we obtain $BIC=1898.5$. Thus, the data clearly favours the Logistic hazard model.
@@ 6.2 Case study II: Breast cancer data
Figure 3 shows the posterior predictive hazard, survival, and “response” functions. Note how the hazard function exhibits an increasing behaviour during the first 2 to 3 years and then decreases and stabilises at a constant value. Interestingly, the equilibrium state is inferred and neither $h$ nor $q$ go to zero. That is, a constant asymptotic hazard $h^{*}$ is predicted for this data set (see Section 3.2). This coincides nicely with previous studies on the evolution of the population hazard function associated with breast cancer patients (Royston and Parmar, 2002). The posterior probability of the equilibrium state may be easily calculated, by checking if $h^{*}>0,q^{*}>0$ and $D>0$ at each iteration of the MCMC. All of our iterations resulted in equilibrium, from an effective sample size of 800. Moreover, the posterior distribution for $h^{*}$ is presented in Figure 3(c). We compare the fit of the hazard-response model against the Weibull and Power Generalised Weibull (PGW) distributions using BIC. The corresponding BICs are $9638.3,9572.0$, for the Weibull and PGW distributions, respectively, and $9561.0$ for the hazard-response model. Thus, the BIC clearly favours the proposed ODE model.
Refer to caption Refer to caption Refer to caption
(a) (b) (c)
Figure 3: rotterdam data: (a) posterior predictive hazard function (red line) and response function (yellow line); (b) survival function (red line) for the Hazard-Response model (7), 0.1 to 0.9 quantiles and the median; and (c) Posterior distribution of $h^{*}$ (left) and $q^{*}$ (right). The posterior predictive for the “response” function $q$ is also included (transparent percentile range) along with the hazard function (a). The system is in equilibrium with probability close to $1$, and $h$ tends to the asymptotic point $h^{*}$.
@@ 7 Discussion
The simulation study presented in this work shows it is possible to recover the true parameter values and the shape of the true hazard functions using the proposed modelling approach. This study also provides guidelines about the effect of the sample size and censoring rates on the inference on the parameters, and the increased uncertainty induced by adding more parameters or additional hidden state variables. The case studies presented here illustrate the use of the proposed modelling approach using real data, as well as the qualitative interpretation of the hazard function in the corresponding context. In particular, Case study II presents an application of the hazard-response model (7), which allows for an interpretation of the competing processes associated to the mortality hazard and the response, at the population level, potentially associated to clinical interventions and the natural immunological response.
A natural extension of our work consists of using other systems of ODEs to model the hazard function. These include the use of other common population growth models (Boyce et al., 2021), or extensions of competition models. For example, similar to Farkas (1984), we could incorporate the value of the hazard function in the past into the effect on the response (i.e. the response depends on the history and evolution of the hazard) as follows:

### v4

@@ 3 Examples of hazard models defined through ODEs
In this section we present three examples that we consider of practical interest. In Section 3.1, we show that the ODE formulation (1) can be used to represent common probability distributions with positive support, but also contains extensions, including distributions with a hazard function $h$ with positive and finite initial condition $h(0)$. A particular model without hidden states, obtained by specifying the hazard function as the solution to the logistic growth model (Boyce et al., 2021), is presented in Section 3.2. This model has been largely used in ecology for modelling population growth, and represents a classical example in most textbooks on ODEs. This model does not include hidden state variables $q$ and has an analytic solution. Section 3.3 introduces a model obtained by defining the hazard function as the solution of a version of the competitive Lotka-Volterra model (Kot, 2001; Murray, 2002). The competitive Lotka-Volterra model is often used to model species in competition, but here we provide an alternative use and interpretation of this model to represent the interaction of the hazard associated with a disease and the response coming from a combination of therapeutic interventions and the immune system (at the population level). This model has a hidden state variable $q$ and no analytic solution, so a numerical solver is employed to approximate its solution. This model also illustrates the richness of solutions one can obtain by adding a hidden state to the logistic growth model.
@@ 3.1 Probability distributions defined through a first order ODE
Note that any given smooth hazard function ${h(t)}$ may be trivially formulated as the solution of an ODE, with $\psi({h},t)={h}^{\prime}(t)$. The following remark shows that our formulation (1) can be used to represent common absolutely continuous probability distributions with positive support by specifying a Bernoulli-type differential equation (Boyce et al., 2021).
@@
The term $a(t)$ in (4) may explicitly depend on $t$, which implies that the corresponding ODE is non-autonomous. If the term $a(t)$ in (4) can be written in terms of $h(t)$ only, then the corresponding ODE is autonomous. Thus, the term $a(t)$ in (4) can be interpreted as an “autonomy coefficient”.
For example, the Weibull hazard function, $h(t)=\beta\kappa t^{\kappa-1}$, may be obtained as the solution to the ODE (4) with $a(t)=(\kappa-1)\left(\dfrac{\beta\kappa}{h(t)}\right)^{\frac{1}{\kappa-1}}-h(t)$, for $\kappa\neq 1$. This is an autonomous ODE. In contrast, for the log-normal distribution $a(t)=\frac{\mu-\sigma^{2}-\log(t)}{\sigma^{2}t}$, which implies that the ODE (4) associated with the log-normal distribution is non-autonomous. This is in line with the classical result that, under mild regularity conditions, autonomous scalar ODEs have monotonic solutions (Wallach, 1948). These ideas are summarised in our context in the following remark.
@@
$h^{\prime}(t)=\psi(h),\quad h(t_{0})=h_{0},$
where $\psi$ is a continuous function, and $h(t)>0$ is the corresponding solution. Then, $h(t)$ is monotonic in $t>0$. Moreover, Let $h(t)>0$ be the solution to a first order (one-dimensional) ODE
$h^{\prime}(t)=\psi(h,t),\quad h(t_{0})=h_{0},$ (5)
If $h(t)$ is monotonic in $t>0$, then the ODE (5) is autonomous. That is, $\psi(h,t)$ depends only on $h$.
@@
Some references have claimed that equations of type (4) can be seen as a characterisation of absolutely continuous distributions with positive support (see Baran and Coughlin, 1987 and Onose, 1986). However, our formulation (1) indicates otherwise, as formulations with hidden state variables cannot be reduced to a Bernoulli ODE of type (4). Moreover, for most of the commonly used distributions, the corresponding hazard function $h(t)$ is the solution to an equation of type (4), with $\lim_{t\to 0}h(t)$ restricted to be either 00 or $\infty$.
3.2 Logistic growth hazard model
Next, we present an example of (1) without hidden state variables, and with positive and finite initial conditions on $h$. The logistic growth hazard model is defined through the following system of ODEs,
@@ 3.2 Logistic growth hazard model
where $\lambda>0$ represents the intrinsic growth rate of the hazard function, $\kappa>0$ represents the upper or lower bound of the hazard function, and $h_{0}>0$ is the value of the hazard function at $t=0$. The logistic growth ODE is an autonomous Bernoulli ODE of type (4), with $a(t)=\lambda-(1+\frac{\lambda}{\kappa})h(t)$. This ODE has the following analytic solution
@@
is a realisation of the logistic growth model. An advantage of model (6) is that it does not force the hazard function to be either 00 or $\infty$ at $t=0$, in contrast to the Weibull distribution (among other distributions). Moreover, the shapes of the hazard function are also increasing or decreasing (see Figure 1). The parameters have a clear interpretation as $h_{0}$ represents the value of the hazard function at $t=0$, $\kappa$ represents the asymptotic point of stability of the hazard function, and $\lambda$ represents the growth rate of the hazard function. Figure 1 shows some examples of the logistic growth hazard model. The numerical solutions are only presented for illustration, since there exists an analytic solution. The numerical solutions were calculated using the “Livermore Solver for Ordinary Differential Equations with Automatic method switching” (LSODA, SciPy implementation). The analytic solution is plotted for $h_{0}=\kappa/10$, for comparison.
In all cases, we can notice how $h(t)\rightarrow\kappa$ as $t\to\infty$.
@@ 3.3 Hazard-Response model
The qualitative analysis of this system is more complicated than that of the simpler logistic-growth model studied in Section 3.2, and may be found in Murray (2002), Section 3.5. Briefly, this model has four steady state solutions. These steady states can be found by solving the equation $\psi_{\bm{\theta}}(h,q)=\mathbf{0}$, which defines four cases: Case 1, $h=0,q=0$; Cases 2 and 3, $h=\kappa,q=0$ and $h=0,q=\kappa$, respectively; and Case 4, $h=h^{*},q=q^{*}$, if both positive, where
@@ 4 Inference
where we are including the initial conditions ${\bf Y}_{0}$ as unknown parameters in this equation. In some cases, as we will discuss in the simulation study and the applications, in Sections 5-6, it is possible to accurately estimate the initial conditions. However, in other cases, as discussed in Section 6.2, the data contains little information about these parameters, leading to practical non-identifiability, which is reflected as flat likelihood surfaces (Cole, 2020). Indeed, the initial conditions are commonly fixed at reasonable values in many practical applications of ODE models (Simpson et al., 2022). Once a solution (analytic or numerical) to the system of ODEs (1) has been obtained, we can evaluate this likelihood function at specific values of the parameters $\bm{\theta}$. In particular, for systems of ODEs without analytic solution and once a solver has been specified, we can retrieve the hazard and cumulative hazard functions, $h(t\mid\bm{\theta},{\bf Y}_{0})$ and $H(t\mid\bm{\theta},{\bf Y}_{0})$, at the sequence of observed time points ${\bf t}=\{t_{1},\dots,t_{n}\}$. This allows for calculating the maximum likelihood estimates (MLEs) of the parameters $\bm{\theta}$ and the initial conditions ${\bf Y}_{0}$ using general-purpose optimisation algorithms (e.g. optim or nlminb in R).
Although it is not possible to come up with general prior choices for the parameters of the hazard models obtained via (1), as different models contain different types of parameters, the ease of interpretation of those parameters facilitates choosing either informative or weakly informative priors. The use of improper priors would require a case-by-case analysis to check the propriety of the posterior distribution.
@@ 5.1 Simulation scenarios
For simulation scenario 1, we simulate $M=250$ samples of sizes $n=250,500,1000,5000$ from the logistic growth hazard model (6) with parameters $(\lambda,\kappa,h_{0})=(0.5,0.05,3.5)$. These parameter values produce a decreasing hazard function, which starts at $h_{0}=3.5$, and decreases at a relatively fast rate, $\lambda=0.5$, to the carrying capacity value $\kappa=0.05$. These samples are simulated using the procedure described in Section 3.2. Administrative censoring times (i.e. fixed) are used to induce censoring rates of approximately $50\%$ and $25\%$.
@@ 5.2 Results
Tables 34 show the summaries of the results for simulation scenario 2. We can see that the Bayes estimates are close to the true values of the parameters, but exhibit some degree of bias which does not disappear with increasing sample size. This seems to be a result of the step size used in the simulation and the interval length. In order to reduce this bias, it is required to use a smaller step size and larger interval length. On the other hand, we can see that this configuration allows one to recover the shape of the hazard with high accuracy, as shown in Figure 7 in the Appendix, which presents the true hazard function together with the $95\%$ predictive intervals.
For benchmarking purposes, we compare the predictive hazard functions obtained in simulation scenario 2 and a B-spline estimator of the hazard function (using the R package bshazard) against the true hazard function. For this purpose, we use the restricted $L_{1}$ distance between two hazard functions, $h$ and $\tilde{h}$, proposed in De Iorio et al. (2009):
[…]
We take $t^{*}$ to be the maximum follow up time for each case. We calculate this distance in two cases: when $\tilde{h}$ represents the predictive hazard function obtained in simulation scenario 2, and when $\tilde{h}$ is the B-spline estimator of the hazard function, with $h_{0}$ being the true hazard function. The corresponding hazard functions are approximated by interpolating the solutions retrieved by the ODE solver and the R package bshazard on a grid of 200 equidistant points on $[0,t^{*}]$ using cubic splines. These approximations are then integrated using the R command integrate. Figures 8-9 in the Appendix display the violin plots corresponding to the $250$ distances obtained in each case. It is evident that the variability of the B-spline estimator consistently results in higher distances from the true generating model.
@@ 6.1 Case study I: Leukemia data
In this application, we analyse the LeukSurv data set from the spBayesSurv R package, using the logistic hazard model (6) presented in Section 3.2. This data set contains information about the survival of $n=1043$ patients with acute myeloid leukemia. The maximum follow-up time was $13.6$ years for this data set. By the end of the follow-up $879$ individuals died and $164$ were still alive.
We simulate from the posterior distribution of the parameters $(\lambda,\kappa,h_{0})$ of the logistic hazard model (6) using the t-walk sampler (Christen and Fox, 2010). The t-walk is a self-adaptive MCMC algorithm that requires no tuning parameters. We obtain $500,000$ posterior samples, and use a burn-in period of $20,000$ iterations and a thinning period of $167$ iterations. This is the estimated integrated auto-correlation time (IAT, see details of output analysis in Molina-Muñoz and Christen, 2022), and divided by the dimension (here, three) results in $55.6$, leading to an effective sample size of $480,000/167\approx 2870$. Figure 2 shows the corresponding posterior predictive hazard and survival functions for this population of cancer patients. The posterior predictive hazard function starts at $h_{0}$ (posterior mean $\approx 3.6$) and sharply decreases to $\kappa$ ($95\%$ posterior credible interval $(0.012,0.125)$) within the first 4 years. These results are in line with the progression of this type of cancer tumours, which tend to develop quickly with only a small portion of the population responding positively to treatments (Shallis et al., 2019), leading to poor prognosis and survival. Now, comparing the logistic hazard model against the Weibull distribution using the Bayesian information criterion (BIC):
$BIC=k\log(n)-2\ell(\lambda,\kappa,h_{0}),$
where $k=3$, the total number of unknown parameters, and $n$ is the sample size. the logistic hazard model has a $BIC=1862.6$, while fitting a Weibull distribution to the same data we obtain $BIC=1898.5$. Thus, the data clearly favours the logistic hazard model.
@@ 6.2 Case study II: Breast cancer data
Figure 3 shows the posterior predictive hazard, survival, and “response” functions, as well as a comparison against a B-spline estimator of the hazard (using the bshazard R package, with 3 degrees of freedom). Note how the hazard function exhibits an increasing behaviour during the first 2 to 3 years and then decreases and stabilises at a constant value. Interestingly, the equilibrium state is inferred and neither $h$ nor $q$ go to zero. That is, a constant asymptotic hazard $h^{*}$ is predicted for this data set (see Section 3.3). This coincides nicely with previous studies on the evolution of the population hazard function associated with breast cancer patients (Royston and Parmar, 2002). The posterior probability of the equilibrium state may be easily calculated, by checking if $h^{*}>0,q^{*}>0$ and $D>0$ at each iteration of the MCMC. All of our iterations resulted in equilibrium, from an effective sample size of 800. Moreover, the posterior distribution for $h^{*}$ is presented in Figure 3(c). Figure 3(d) illustrates the large variability of the B-splines estimator of the hazard function compared to the hazard-response model. The “wiggliness” of the B-splines estimator of the hazard function complicates making precise epidemiological statements about the evolution of the disease. Finally, we compare the fit of the hazard-response model against the Weibull and Power Generalised Weibull (PGW, 3 parameters) distributions using BIC
$BIC=k\log(n)-2\ell(\lambda,\kappa,\alpha,\beta),$
where $k=4$, the total number of unknown parameters, and $n$ is the sample size. The corresponding BICs are $9638.3,9572.0$, for the Weibull and PGW distributions, respectively, and $9561.0$ for the hazard-response model. Thus, the BIC clearly favours the proposed ODE model.
Refer to caption Refer to caption
(a) (b)
Refer to caption Refer to caption
Figure 3: rotterdam data: (a) posterior predictive hazard function (red line) and response function (yellow line); (b) survival function (red line) for the Hazard-Response model (7), 0.1 to 0.9 quantiles and the median; (c) Posterior distribution of $h^{*}$ (left) and $q^{*}$ (right); and (d) posterior predictive hazard function (black line) and B-spline estimator (gray line) with $95\%$ confidence interval. The posterior predictive for the “response” function $q$ is also included (transparent percentile range) along with the hazard function (a). The system is in equilibrium with probability close to $1$, and $h$ tends to the asymptotic point $h^{*}$.
@@ 7 Discussion
The simulation study presented in this work shows it is possible to recover the true parameter values and the shape of the true hazard functions using the proposed modelling approach. This study also provides guidelines about the effect of the sample size and censoring rates on the inference on the parameters, and the increased uncertainty induced by adding more parameters or additional hidden state variables. The case studies presented here illustrate the use of the proposed modelling approach using real data, as well as the qualitative interpretation of the hazard function in the corresponding context. In particular, Case study II presents an application of the hazard-response model (7), which allows for an interpretation of the competing processes associated to the mortality hazard and the response, at the population level, potentially associated to clinical interventions and the natural immunological response. We conducted model selection using the classical definition of the Bayesian Information Criterion (BIC). Alternatively, one could employ modified versions of BIC tailored for censored samples, or opt for Bayesian methods for model selection, such as the use of Bayes factors or model posterior probabilities.
A natural extension of our work consists of using other systems of ODEs to model the hazard function. These include the use of other common population growth models (Simpson et al., 2022), such as Gompertz or Richard’s models, or extensions of competition models. For instance, similar to Farkas (1984), we could incorporate the value of the hazard function in the past into the effect on the response (i.e. the response depends on the history and evolution of the hazard) as follows:
@@
$25\%$ censoring
Refer to caption Refer to caption
(a) $n=250$ (b) $n=500$
Refer to caption Refer to caption
(c) $n=1000$ (d) $n=5000$
Figure 8: Simulation scenario 2, $25\%$ censoring: Distance between the posterior predictive hazards and the B-Spline estimator to the true generating model.
$50\%$ censoring
Refer to caption Refer to caption
(a) $n=250$ (b) $n=500$
Refer to caption Refer to caption
(c) $n=1000$ (d) $n=5000$
Figure 9: Simulation scenario 2, $50\%$ censoring: Distance between the posterior predictive hazards and the B-Spline estimator to the true generating model.